# A low complexity contextual stacked ensemble-learning approach for pedestrian intent prediction

### v1

@@ Abstract
Walking as a form of active travel is essential in promoting sustainable transport. It is thus crucial to accurately predict pedestrian crossing intention and avoid collisions, especially with the advent of autonomous and advanced driver-assisted vehicles. Current research leverages computer vision and machine learning advances to predict near-misses; however, this often requires high computation power to yield reliable results. In contrast, this work proposes a low-complexity ensemble-learning approach that employs contextual data for predicting the pedestrian’s intent for crossing. The pedestrian is first detected, and their image is then compressed using skeleton-ization, and contextual information is added into a stacked ensemble-learning approach. Our experiments on different datasets achieve similar pedestrian intent prediction performance as the state-of-the-art approaches with $99.7\%$ reduction in computational complexity. Our source code and trained models will be released upon paper acceptance
@@ I Introduction
Active mobility is a critical component in the pathway to sustainable and effective urban transportation systems. Active travel is a transportation mode in which people travel from one place to another by being physically active, such as walking or cycling. However, the growing concerns about the safety of vulnerable road users (VRUs), such as pedestrians, have slowed the uptake of active travel. According to the World Health Organisation data, $1.19$ million people die each year as a result of road traffic crashes and more than half are VRUs [^1] https://www.who.int/news-room/fact-sheets/detail/road-traffic-injuries.
@@
The near future vision of urban mobility consists of various modes of transport sharing the open urban space and road network including VRUs, driver-assisted, and driverless vehicles [1]. Human drivers have ‘common sense’ that allows them to ‘read’ a scene, predict the behavior of VRUs, and communicate with them through eye contact to indicate safety to crossing [2]. Can we train autonomous vehicles to develop a similar ‘common sense’ that predicts the behavior of VRUs? Pedestrian intent prediction (PIP) has gained increased attention in recent years, with significant progress being made in autonomous vehicles as a part of the intelligent transportation systems paradigm.
@@
PIP methods often build on technologies for connected and autonomous vehicles (CAVs) and leverage a spectrum of sensors (e.g., radar, LIDAR, optical cameras, and ultrasonic sensors) with different machine learning (ML) methods to automate decisions that avoid collisions [3]. Among these sensors, cameras provide the most detailed depiction of the vehicle’s surroundings, and the advances in related computer vision approaches allow for precise localization of pedestrians and PIP. However, camera-based data is often complex, requiring high computation power for processing. Furthermore, this data carries personal information about pedestrians that might put them at risk if misused while being irrelevant in solving the PIP problem.
@@
There are two dominant approaches for computer vision (CV)-enabled PIP: the first is based on cameras installed along the roads, and the second relies on car-mounted cameras. The former relies on fixed cameras; thus, they require reliable broadband connectivity to the cloud server, which has high computational power and storage [4]. Other work leverages edge computing to offload pedestrian-related prediction tasks and ensure reliability [5]. Privacy concerns arise when data containing personal information (e.g., people’s faces) is sent over the network for processing. The second dominant approach for enabling PIP is based on collecting and processing data from vehicle-mounted cameras and other sensory devices that are key enablers for developing the advanced driver-assistance system (ADAS) and automated driving technologies. With the advent of 5G and next-generation networks, broadband connectivity is possible for moving vehicles but is less reliable and more costly than the connectivity of fixed cameras [6]. The processing power available to ADAS is less limited than road-based cameras, but the robustness of the sensors, data fusion, and AI algorithms varies significantly between different vendors. It follows that centralized learning is still preferred to ensure the reliability of the AI models. Therefore, data communication to the central server is predominant even in this approach and suffers from the challenge of transferring large volumes of data [7].
@@
Another rising concern common to both approaches stems from their heavy dependency on computationally complex AI models and the contribution of these to greenhouse gas emissions. This is underscored in [8] in which computational researchers are solicited to monitor and report the carbon footprint of their work to ensure a more sustainable future. The relation between computational complexity and carbon footprint is highlighted in [9] which acknowledges the irrefutable role of AI in enabling sustainable futures while it advocates for effective low-complexity AI models.
@@
Recent research [10] [11] [7] proposes Graph Neural Networks to model the change of human keypoints in time to represent pedestrian crossing behavior. Others [12] employ Long Short-Term Memory (LSTM) to model the relationship between pedestrian locations and crossing behavior. A transformer network is proposed in [13] for PIP to capture the AI uncertainty in scene complexities. Despite the incremental improvements in PIP performance in these works, it often comes at the cost of higher computational complexity and greater impact on greenhouse gas emissions.
@@
This work proposes a new Contextual Stacked Ensemble-learning (CSE) method for solving PIP using car-based cameras. Similar to [10] [11] [7], we use skeleton-ization, which compresses the image of a pedestrian to 17 keypoints, thus reducing the volume of the camera-based data by $4336$ times (see Section IV-B1) while addressing privacy concerns. As an outcome, transmitting the skeleton-ized data to the cloud server for further processing and conducting a deep intent prediction becomes feasible. The main goals of the proposed method are threefold: data compression and privacy, robust and reliable PIP, and low computational complexity. The robust and reliable PIP model is achieved by leveraging contextual information (e.g., categorical data describing the scene and pedestrian trajectory) with stacked ensemble learning. The proposed CSE approach achieves comparable PIP performance to state-of-the-art work with a step reduction in complexity ($99.99\%$ reduction in FLOPS and $99.7\%$ less trainable parameters). Our approach received the first prize award of the IEEE Intelligent Transportation Systems Society (ITSS) Student Competition in Pedestrian Behavior Prediction [14].
@@ II Problem formulation
In this work, the PIP problem is formulated as predicting the behavior of a pedestrian as Crossing or Not-Crossing based on video clips taken from the front of a vehicle. Each video clip, denoted as $\mathbb{S}_{v}$, for $v\in\{1,\dots,V\}$ ($V$ is the total number of clips in the dataset), contains a consecutive set of $M$ frames $F_{i}$, where $i$ is the index of frames $i\in\{1,\dots,M\}$. Each clip $\mathbb{S}_{v}$ includes a maximum of $P$ bounding boxes (bbx), and each bbx frames a single pedestrian with an index $j$ and may include other pedestrians in the background (i.e., not labeled). It follows that each annotated clip includes the labeled behavior of $P$ pedestrians with a unique index $j\in\{1,\dots,P\}$. Given the one-to-one mapping between bbx and pedestrians, the same index $j$ is used to reference the bbx that frames a pedestrian with index $j$. In each video clip $\mathbb{S}_{v}$, a pedestrian with index $j\in\{1,\dots,P\}$ (and the bbx that frames them) is annotated with a label $y$; $y=1$ Crossing if pedestrian intends to cross a street at any point in $\mathbb{S}_{v}$ or $y=0$ for (Not-Crossing or Irrelevant, i.e., without a behavioral label) otherwise.
The data in each video clip is analysed from two perspectives. The first relates to the pedestrian and is based on sliding widows of $32$ frames that span the full video clip. Two data streams are extracted from this perspective: keypoints that represent the pose of the pedestrian (see Section IV-B1) and the pedestrian’s trajectory relative to the frame (see Section IV-B3). The other perspective captures the contextual information from two angles: global context and local context. The former is common to all pedestrians in a given frame and describes the speed of the ego-vehicle, the status of the traffic light it faces, and the road type it drives on. The latter is specific to each pedestrian and describes their relative local context such as standing at an intersection (see Section IV-B2). The goal of this work is therefore to classify sliding windows as positive (Crossing) or negative (Not-Crossing and Irrelevant) case.
@@ III Methodology
Different methods have been proposed to address PIP in [15, 7][16, 17] however, these methods fail to examine class performance or suffer from high computational complexity. Therefore, in this work, we propose Contextual Stacked Ensemble (CSE), a novel ensemble learning system that builds on three individual models which we introduce in this section. The first model, M1, consists of a graph neural network (GNN) model for classifying skeleton-ized data, $\mathbb{K}_{v}$ as Crossing/Not-Crossing. The second model, M2, employs stacked gated recurrent network units (GRUs) for solving PIP using time-series and categorical data from JAAD’s vehicle and traffic annotations. The third model, M3, employs a one-dimensional convolutional neural network for solving PIP based on a pedestrian’s trajectory, $\boldsymbol{\tau_{v}}$. The CSE combines these three models as shown in Figure 5 (right). For comparison, two models are also implemented using a similar structure. The entire ensemble learning system is thus structured as Figure 1 with detailed explanations of each of its three models (M1, M2, M3) in the following subsections.
@@ III-A M1: Graph Neural Network
The GNN model (M1) is used to classify a sequence of graphs in which the vertices describe the $K_{max}$ keypoints in the frame with index $i$ and are represented by the matrix $\mathcal{K}_{i}$ which is obtained through the skeleton-ization process; in this case, $k_{\kappa}=\{w_{\kappa},h_{\kappa}\}$ represents the relative position of a keypoint with index $\kappa$ and $\mathbb{K}_{v}$ represents the sequence of matrices $\mathcal{K}_{i}$ during the video clip $\mathbb{S}_{v}$.
@@
The model consists of a convolutional block, a skip connection and a GRU, as shown in Figure 2. The input keypoints are first applied matrix multiplication with a normalized adjacency matrix. The matrix product involves summing up keypoint features from adjacent nodes. The summed-up features of each node are embedded into a convolutional layer for further feature extraction. To prevent shrinking feature maps, the convolutional layer must use “same” padding. A dropout (0.5) layer is then used to encourage model generalization.
The skip connection operation adds input keypoints to the dropout output. By passing coming layers and preserving early layer information, the model can keep both low and high level representations, thus improves feature learning. A 1x1 convolution on the skip connection branch adjusts the channel dimension to enable the addition to the features output by the 2D convolutional block.
Finally, a GRU layer is used to model the evolution of keypoint features in the time domain. The hidden size is set to $16$. The last hidden output is then fed to a fully-connected layer for Crossing/Not-Crossing classification.
Figure 2: M1: The figure shows model structure with input and output dimension in each model block.
@@
Inspired by the stacked GRU in [18], we apply the stacked GRU in modelling time-series and categorical contextual information from the vehicle and traffic background as shown in Figure 3. The stacked GRUs model (M2) iterates each time-series data and feeds it in a GRU with only one layer to produce all hidden outputs from frame $F_{1}$ to $F_{max}$. Before feeding all hidden outputs to the next GRU, a next time-series data is merged with the hidden outputs. The output of the last GRU is the last hidden output instead of a full sequence of hidden outputs.
@@
  • Data 4- (categorical with one value for each video)- Type of road {Garage, Parking lot, Street}: It is more likely that a pedestrian seen on the street might cross than a pedestrian standing/walking in a parking lot.
@@
Figure 3: M2: a stacked GRUs is used to model contextual information that can affect crossing decision.
@@ III-B M3: 1D convolutional Neural Network
For modelling the relationship between pedestrian trajectory $\boldsymbol{\tau_{v}}$ and PIP decision, a shallow 1D-CNN (M3) is proposed (see Figure 4). 1D-CNN is often used to predict a future trajectory given historical data[19, 20] while [21] uses both RNN and CNN to predict future classes of an encoded trajectory. Trajectories possess not only pedestrian speed, direction, and pedestrian depth information but are also strongly related to PIP decisions. We, therefore, adopt a 1D-CNN for M3 to solve PIP based on past trajectory as contextual information. Our proposed 1D-CNN model contains only three layers: a convolutional layer for extracting 1D time-series features, a dropout layer for preventing overfitting, and a fully connected layer for PIP classification, as shown in Figure 4.
Figure 4: M3: A 1D-CNN could be used to predict future trajectory given historical data. The five inputs of the trajectory are introduced in Section IV-B3.

### v2

@@ Abstract
Walking, as a form of active travel, is a fundamental element in promoting sustainable transport. Accurately predicting pedestrian crossing intention is vital in preventing collisions, especially with the rise of autonomous vehicles and the Internet-of-Vehicles (IoV). Current research relies on computer vision and machine learning to predict near-misses, often requiring high computation power to yield reliable results. In contrast, our work presents a low-complexity ensemble-learning method that utilizes contextual data to predict pedestrian intent for crossing. The pedestrian is first detected, and their image is compressed using skeleton-ization. Next, we incorporate contextual information into a stacked ensemble-learning model. Our experiments conducted on different datasets show that our approach achieves a similar pedestrian intent prediction performance to the state-of-the-art approaches, with a $99.97\%$ reduction in computational complexity. Our source code and trained models will be released upon paper acceptance.
@@ I Introduction
Active travel is critical in the pathway to sustainable and effective urban transportation systems. Active travel is a transportation mode in which people travel from one place to another by being physically active, such as walking or cycling. However, the growing concerns about the safety of vulnerable road users (VRUs), such as pedestrians, have slowed the uptake of active travel. According to the World Health Organization data, $1.19$ million people die each year as a result of road traffic crashes, and more than half are VRUs [^1] https://www.who.int/news-room/fact-sheets/detail/road-traffic-injuries.
@@
The future vision of urban mobility consists of various modes of transport sharing the open urban space and road network, including VRUs, connected and autonomous vehicles as part of the Internet of Vehicles (IoV) [1]. Human drivers have ‘common sense’ that allows them to ‘read’ a scene, predict the behavior of VRUs, and communicate with them through eye contact to indicate safety to crossing [2]. Can we train autonomous vehicles to develop a similar ‘common sense’ that predicts the behavior of VRUs? Pedestrian intent prediction (PIP) has gained attention recently, particularly due to advancements in autonomous vehicles within the intelligent transportation systems framework.
@@
PIP methods often rely on technologies used in connected and autonomous vehicles (CAVs) and utilize a variety of sensors, including radar, LIDAR, optical cameras, and ultrasonic sensors. These sensors leverage various machine learning (ML) techniques to automate decision-making process that helps avoid collisions [3]. Recent advances in computer vision have greatly improved the ability to identify pedestrians and PIPs accurately. However, the data collected by cameras is often complex and requires high computation power for processing. Furthermore, this data may contain personal information about pedestrians, which could pose a risk if misused while being irrelevant to addressing the PIP problem.
@@
There are two primary approaches for computer vision (CV)-enabled PIP. The first approach relies on cameras installed on the roads, which require reliable broadband connectivity to a cloud server that offers high computational power [4]. However, this cloud processing can raise privacy concerns, mainly when personal information (e.g., people’s faces) is transmitted over the network for analysis. Some research addresses this issue using edge computing to offload pedestrian-related prediction tasks and improve reliability. The second approach involves collecting and processing data from cameras and other sensory devices mounted on vehicles. This approach is essential for developing advanced driver-assistance systems (ADAS) and IoV. Integrating ADAS with IoV offers opportunities to improve safety, decision-making, and efficiency in driving.
With the emergence of 5G and next-generation networks, broadband connectivity for moving vehicles is possible. However, this connectivity tends to be less reliable and more costly compared to the connectivity of fixed cameras [5]. Moreover, the robustness of the sensors and of the data fusion and artificial intelligence (AI) algorithms varies significantly among different vendors. As a result, centralized learning is often preferred to ensure the reliability of AI-based decisions. Therefore, data communication to the central server remains predominant in this approach, but it faces challenges due to the large volumes of data that need to be transferred [6].
@@
A growing concern shared by both approaches is their significant overhead resulting in computationally complex AI models, which contribute to greenhouse gas emissions. This issue is emphasized in a study where computational researchers are encouraged to monitor and report the carbon footprint of their work to promote a more sustainable future [7]. The relation between computational complexity and carbon footprint is highlighted in [8], where the study acknowledges the essential role of AI in facilitating sustainable outcomes while advocating for the development of efficient and low-complexity AI models.
This work proposes a new method called Contextual Stacked Ensemble-learning (CSE) for solving PIP based on data captured by front car-mounted cameras. Similar to previous studies [9], [10], [6], we employ a process known as skeleton-ization, which compresses the image of a pedestrian into 17 keypoints. This approach reduces the volume of camera-based data by 4,336 times (see Section IV-B1), while also addressing privacy concerns. As a result, it becomes feasible to transmit the skeletonized data to a cloud server for further processing and to perform deep intent prediction.
@@
In recent research, Graph Neural Networks are proposed to model the change of human keypoints in time, which helps represent pedestrian crossing behavior[9] [10] [6]. Other studies [11] employ Long Short-Term Memory (LSTM) to model the relationship between pedestrian locations and crossing behavior. A transformer network has also been introduced in [12] for PIP to capture the AI uncertainty in scene complexities. This research has advanced the PIP performance but such methods often require high computational complexity and have a significant impact on greenhouse gas emissions.
@@
To this end, the main objectives of our proposed method are to ensure robust and reliable PIP with low computational complexity and data compression to enhance privacy. The robust and reliable PIP model is achieved by leveraging contextual information, such as categorical data regarding the scene and pedestrian trajectory, through stacked ensemble learning. The proposed CSE approach achieves comparable PIP performance to state-of-the-solutions while significantly reducing complexity, achieving ($99.99\%$ reduction in FLOPS and $99.97\%$ less trainable parameters). Moreover, our approach won the first prize award of the IEEE Intelligent Transportation Systems Society (ITSS) Student Competition in Pedestrian Behavior Prediction [13].
@@ II Problem formulation
In this work, the PIP problem is formulated as predicting the behavior of a pedestrian between Crossing or Not-Crossing based on video clips taken from a camera mounted at the front of a vehicle. Each video clip, denoted as $\mathbb{S}_{v}$, for $v\in\{1,\dots,V\}$ ($V$ is the total number of clips in the dataset), contains a consecutive set of $M_{v}$ frames denoted as $F_{i}$, where $i$ is the index of frames $i\in\{1,\dots,M_{v}\}$. Each clip $\mathbb{S}_{v}$ includes a maximum of $P_{v}$ bounding boxes (bbx), and each bbx bounds a single pedestrian with an index $j$ and may include other pedestrians in the background (i.e., not labeled). We note that video clips differ in length $M_{v}$ and include a different number $P_{v}$ of captured pedestrians. It follows that each annotated clip includes the labeled behavior of $P_{v}$ pedestrians with a unique index $j\in\{1,\dots,P_{v}\}$. Given the one-to-one mapping between bbx and labeled pedestrians, the same index $j$ is used to reference the bbx that bounds a pedestrian with index $j$.
A sliding windows of width $F_{max}=32$ frames is used to span each full video clip to generate segments with index $s\in\{1,\dots,M_{v}/F_{max}\}$ that describe the movement of pedestrian with index $j$ in a video clip with index $v$.
In each video clip $\mathbb{S}_{v}$, a pedestrian with index $j\in\{1,\dots,P_{v}\}$ (and the bbx that frames them) is annotated with a label $y$; $y=1$ Crossing if pedestrian intends to cross a street at any point in $\mathbb{S}_{v}$ or $y=0$ for (Not-Crossing or Irrelevant, i.e., without a behavioral label), otherwise.
The PIP problem can thus be formulated as a classification predictive modeling with the objective of predicting the label $\hat{y}_{v,i}$ associated with each segment $p_{v,i,s}$ of a pedestrian bounded by bbx with index $j$, in sliding window $s$ of video clip $\mathbb{S}_{v}$. It follows that this problem is concerned with maximizing the number of correctly predicted labels $\hat{y}_{v,j,s}$} for all segments $p_{v,j,s}\in\mathbb{P}$ and $\forall v\in\{1,\dots,V\},\forall j\in\{1,\dots,P_{v}\}$, and $\forall s\in\{1,\dots,M_{v}/F_{max}\}$, as formulated here:
$\max~~~~~\sum_{v=1}^{V}\sum_{j=1}^{P_{v}}\sum_{s=1}^{\frac{M}{F_{max}}}(\hat{y}_{v,j,s}=y_{v,j,s})$ (1)
Towards solving the problem in (1), we need to first identify the features that describe the segments in $\mathbb{P}$. To this end, the data in each video clip are analyzed from two perspectives. The first examines the movement of target pedestrians by studying the corresponding bbx and its changing relative position to the frame. Two data streams are extracted from this perspective: keypoints that represent the pose of the pedestrian (see Section IV-B1) and the pedestrian’s trajectory relative to the frame (see Section IV-B3). The other perspective captures the contextual information from two angles: global context and local context. The former is common to all pedestrians in a given frame and describes the speed of the ego-vehicle, the status of the traffic light it faces, and the road type it drives on. The latter is specific to each pedestrian and describes their relative local context such as standing at an intersection (see Section IV-B2). The goal of this work is therefore to first generate these features that will then be used to classify segments in $\mathbb{P}$ as positive (Crossing) or negative (Not-Crossing and Irrelevant) case.
@@ III Methodology
Different methods have been proposed to address PIP in [14, 6][15, 16] however, these methods suffer from high computational complexity. Therefore, in this work, we propose Contextual Stacked Ensemble (CSE), a novel ensemble learning system that builds on three individual models which we introduce in this section. The first model, M1, consists of a graph neural network (GNN) model for classifying skeleton-ized data, $\mathbb{K}_{v}$ as Crossing/Not-Crossing. The second model, M2, employs stacked gated recurrent network units (GRUs) for solving PIP using time-series and categorical data from JAAD’s vehicle and traffic annotations. The third model, M3, employs a one-dimensional convolutional neural network (1D-CNN) for solving PIP based on a pedestrian’s trajectory, $\boldsymbol{\tau_{v}}$. The CSE combines these two or three models as shown in Figure 5 (right). Each of models M1, M2, M3, and CSE model are structured to represent the PIP of a target pedestrian during a sliding window of $F_{max}$ frames. For comparison, two models are also implemented using a similar structure. The entire ensemble learning system is thus structured as shown in Figure 1.
@@ III-A M1: Graph Neural Network
The GNN model (M1) is used to classify a sequence of graphs where vertices are represented by the $K_{max}$ keypoints representing a target pedestrian with index $j\in\{1,\dots,P_{v}\}$ bounded by a bbx in the sliding window with index $s\in\{1,\dots,M_{v}/F_{max}\}$ of clip with index $v$ and maximum number of indexed pedestrians $P_{v}$. For simplicity, in the description of M1, we drop the indices related to the video clip and to the pedestrian and focus on the index that tracks the movement of the target pedestrian, i.e., the index of the frame. For a given target pedestrian bounded by a bbx, each vertex (i.e., keypoint) in each frame $F_{i}$ within the sliding window is depicted by a matrix $\mathcal{K}_{i}$, which is obtained through the skeleton-ization process. In this context, $k_{\kappa}=\{w_{\kappa},h_{\kappa}\}$ denotes the relative position of a keypoint with index $\kappa$, and $\mathbb{K}_{v}$ (see Figure 1) represents the sequence of matrices $\mathcal{K}_{i}$ throughout a video clip $\mathbb{S}_{v}$.
@@
The model consists of a Convolutional Graph Neural Network layer (GCN layer), a Relu+Dropout, a skip connection, a Flatten and a GRU, as shown in Figure 2. The input keypoints are first applied matrix multiplication with a normalized adjacency matrix. The matrix product involves summing up keypoint features from adjacent nodes. The summed-up features of each node are embedded into a convolutional operation in the GCN layer for further feature extraction. To prevent shrinking feature maps, the convolutionaloperation must use “same” padding. A Relu and Dropout (0.5) layers are then used to encourage model generalization.
The skip connection operation adds input keypoints to the dropout output. By passing coming layers and preserving early layer information, the model can keep both low and high level representations, thus improves feature learning. A 1x1 convolution on the skip connection branch adjusts the channel dimension to enable the addition to the features output by the GCN layer.
Finally, after flattening the merged features from Dropout and skip connection, a GRU layer is used to model the evolution of keypoint features in the time domain. The hidden size is set to $16$. The last hidden output is then fed to a fully-connected layer for Crossing/Not-Crossing classification.
Figure 2: M1: The figure shows the model structure with input and output dimensions in each model block. Note that the shape of Keypoints at the input of M1 is determined by $F_{max}=32$, 2D coordinates, and $K_{max}=14$.
@@
Inspired by the stacked GRU in [17], we apply the stacked GRU in modelling time-series and categorical contextual information from the vehicle and traffic background as shown in Figure 3. The stacked GRUs model (M2) iterates each time-series data and feeds it to a GRU with only one layer to produce all hidden outputs from frame $F_{1}$ to $F_{max}$. Before feeding all hidden outputs to the next GRU, a next time-series data is merged with the hidden outputs. The output of the last GRU is the last hidden output instead of a full sequence of hidden outputs.
@@
  • Data 4 (categorical with one value for each video)- Type of road {Garage, Parking lot, Street}: It is more likely that a pedestrian seen on the street might cross than a pedestrian standing/walking in a parking lot.
@@
Figure 3: M2: a stacked GRUs is used to model contextual information that can affect the crossing decision. Note: data 1 is the vehicle speed, data 2 is the traffic signal status, data 3 is the pedestrian-intersection status, and data 4 is the type of road. The shape of each input data corresponds to the size of the sliding window $F_{max}$.
@@ III-B M3: 1D convolutional Neural Network
For modelling the relationship between pedestrian trajectory $\boldsymbol{\tau_{v}}$ and PIP decision, a shallow 1D-CNN (M3) is proposed (see Section 4). 1D-CNN is often used to predict a future trajectory given historical data[18, 19] while [20] uses both RNN and CNN to predict future classes of an encoded trajectory. Trajectories possess not only pedestrian speed, direction, and pedestrian depth information but are also strongly related to PIP decisions. We, therefore, adopt a 1D-CNN for M3 to solve PIP based on past trajectory as contextual information. Our proposed 1D-CNN model contains only three layers: a convolutional layer for extracting 1D time-series features, a dropout layer for preventing overfitting, and a fully connected layer for PIP classification, as shown in Figure 4.
Figure 4: M3: A 1D-CNN could be used to predict future trajectory given historical data. The five inputs of the trajectory are introduced in Section IV-B3. Note that the input dimension of the first Conv1D is ($F_{max}$, five trajectory attributes) = $(32,5)$. They are normalized bounding box height, width and size, the change of size and the speed of the box.